The age of a man is 10 years more than three times his brother's age. Find the age of each if 5 years from now he will be 4 years more than twice his brother's age.
Answer by ankor@dixie-net.com(22740)

The age of a man is 10 years more than three times his brother's age.
Find the age of each if 5 years from now he will be 4 years more than twice his brother's age.
:
Let m = man's present age
Let b = brother's present age
:
Write an equation for each statement.
:
"The age of a man is 10 years more than three times his brother's age."
m = 3b + 10
:
"5 years from now he will be 4 years more than twice his brother's age."
m + 5 = 2(b+5) + 4
m + 5 = 2b + 10 + 4
m = 2b + 14 - 5
m = 2b + 9
From the 1st equation, replace m with (3b+10)
3b + 10 = 2b + 9
3b - 2b = 9 - 10
b = -1, obviously, this is a nonsense problem
:
m = 3(-1) + 10
man 7 yrs old, brother is -1 (Won't be born for a year, I guess)